# Seo Hyun Kim1     Kai Tzu-iunn Ong1 Taeyoon Kwon1     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1

### v1

@@ Abstract
Large language models (LLMs) are capable of processing lengthy dialogue histories during prolonged interaction with users without additional memory modules; however, their responses tend to overlook or incorrectly recall information from the past. In this paper, we revisit memory-augmented response generation in the era of LLMs. While prior work focuses on getting rid of outdated memories, we argue that such memories can provide contextual cues that help dialogue systems understand the development of past events and, therefore, benefit response generation. We present [Uncaptioned image]Theanine, a framework that augments LLMs’ response generation with memory timelines – series of memories that demonstrate the development and causality of relevant past events. Along with Theanine, we introduce TeaFarm, a counterfactual-driven question-answering pipeline addressing the limitation of G-Eval in long-term conversations. Supplementary videos of our methods and the [Uncaptioned image]TeaBag dataset for TeaFarm evaluation are in https://theanine-693b0.web.app/.
[Uncaptioned image]Theanine: Revisiting Memory Management in Long-term Conversations with Timeline-augmented Response Generation
Seo Hyun Kim1thanks: These two authors are the co-first authors and both have the right to list their names first on their CVs.     Kai Tzu-iunn Ong1 Taeyoon Kwon1thanks: T Kwon, N Kim, and K Ka are the co-second authors.     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1 1Yonsei University, 2Seoul National University
@@ 1 Introduction
Recently, the impressive performance of large language models (LLMs) has caused a trend of having longer conversations with them across multiple sessions (Fui-Hoon Nah et al., 2023). Such prolonged interactions require the models to memorize previous events or speaker information, tailoring the responses accordingly to keep up user interest (Adiwardana et al., 2020). While the large context windows of LLMs [^1] Recent LLMs, such as GPT-4 (Achiam et al., 2023) and LLama3 (MetaAI, 2024), have contexts windows of 128,000 and 1,040,000 tokens, respectively. allow them to process the entire dialogue history for response generation (RG), this often leads to biased attention toward the latest user utterance, ignoring relevant contexts from the past (Liu et al., 2024), as presented in Figure 1 (a).
Figure 1: Empirical examples of failed responses due to (a) bias to the latest input (b) the absence of an important past event (“afraid of cruise ships”) on the timeline. (c) shows a response augmented with the memory timeline.
@@
One straightforward solution is to condense past conversations into summarized memories and retrieve them to augment response generation in later sessions (Xu et al., 2022a, b; Lu et al., 2023), but the growing span of memories can hinder the quality of memory retrieval as dialogue sessions accumulate. Although this can be mitigated by updating the older memories (Bae et al., 2022; Zhong et al., 2024; Kim et al., 2024a), we observe severe information loss. As illustrated in Figure 1 (b), an earlier memory on the timeline, which contains an important speaker persona (i.e., “afraid of ships”), is removed during memory update (thus can not be retrieved), yielding a suboptimal response.
Motivated by these, in this paper, we revisit memory-augmented response generation for long-term conversations and address the aforementioned bottlenecks. Specifically, we focus on using the timeline built with relevant memories to augment response generation. To this end, we present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019). a framework of timeline-augmented chain-of-thought reasoning for response generation in long-term conversations.
Inspired by how humans naturally link new memories to existing memories of relevant events and retrieve them as a whole (Bartlett, 1995), Theanine manages memories with a graph structure, i.e., the memory graph, where memories are linked based on how they relate to each other (Phase I). Intuitively, this can be done by linking each memory to others that have high text similarity. Yet, we find this naive approach not much beneficial to response generation, due to the plain connection between memories (e.g., “this happened $\rightarrow$ that occurred”). To resolve this, we leverage an LLM to dynamically link memories based on their temporal and cause-effect commonsense relations (Hwang et al., 2021) via our designed algorithm. Next, in memory retrieval for generating a new response (Phase II-1), instead of solely retrieving top-$k$ memories using current conversation as query, we retrieve the whole memory timeline representing the development of relevant events. Then, to tackle the discrepancy between off-line memory construction and online deployment (Phase II-2), Theanine refines the retrieved timeline such that it provides tailored information for the current conversation. Finally, in response generation (Phase III), we leverage LLMs’ chain-of-thought (CoT) reasoning ability (Wei et al., 2022) to reason over current conversation and refined timelines, conclude useful information, and generate the final response. Our contributions are three-fold:
  • We revisit memory-augmented response generation in long-term conversations with a novel timeline-augmented framework, [Uncaptioned image]Theanine. In both LLM-based (i.e., G-Eval) and human evaluation, Theanine demonstrates (i) better responses that are detailed/non-generic and (ii) efficiency in properly referencing past conversations.
@@
  • The lack of ground-truth mapping between ongoing conversations (queries) and memories for retrieval poses a challenge in evaluating memory-augmented dialogue systems. We present TeaFarm, a counterfactual-driven question-answering pipeline, that assesses systems in referencing past conversations, which does not require human intervention.
  • We further curate [Uncaptioned image]TeaBag, the counterfactural question-answer data for TeaFarm evaluation on two widely used benchmark datasets of long-term conversations: Multi-Session Chat (MSC) (Xu et al., 2022a) and Conversation Chronicles (CC) (Jang et al., 2023).
@@ 2 Methodologies
We present [Uncaptioned image]Theanine (Figure 2), a framework that resembles the mechanism of how humans store and retrieve memories for conversations:
@@ 2.1 Memory Graph Construction (Phase I)
To replicate how humans construct memories, we approach memory management based on a graph structure, i.e., the memory graph $G$:
@@
Humans, on the other hand, interpret events by considering the relation between them, such as “how does an event affect the other?” or “why did this person make that change?”.
Therefore, we adopt a relation-aware memory linking, where each edge between two memories is encoded with their cause-effect commonsense relation $r\in R$, along with their temporal order. These relations are adopted from Hwang et al. (2021) and modified by us, including HinderedBy, Cause, Want, and 4 more (Appendix A.1).
@@
where $\Upsilon(\cdot,m_{new})\in R$ indicates that the given memory is assigned with an $r\in R$ with $m_{new}$, [^4] Limited by the performance of retrievers, it is possible that an $m\in M_{a}$ does not have a relation, other than just textual overlap, with $m_{new}$. We address this by allowing the LLM to output “None”. and such assigned memories are defined as $M_{a}^{*}$.
@@
We then proceed to link $m_{new}$ to the graph, a straightforward approach is linking $m_{new}$ to all $m\in M_{a}^{*}$. Yet, we find it algorithmically inefficient (Figure 7). Thus, we selectively link $m_{new}$ to only those $m\in M_{a}^{*}$ that are recent. To identify those recent memories, we first locate every connected component $C_{i}\subset G^{t}$ that contains at least one $m\in M_{a}^{*}$, as shown in Figure 3 (a) and (b):
where $\mathbb{C}$ is the collection of those $C$ and $\mathtt{V}(\cdot)$ represents “vertices in”. Now, we link $m_{new}$ to the most recent $m\in M_{a}^{*}$ in each $C_{i}\subset\mathbb{C}$ (Figure 3 (c)). Therefore, the memories $M_{linked}$ that are linked to $m_{new}$ is defined as follows:
where $\Omega(\cdot)$ indicates “the most recent memory in”. Linking all memories from session $t$ to $G^{t}$, we then obtain a new memory graph $G^{t+1}$. The pseudo algorithm for Phase I is in Algorithm 1.
@@ 2.2 Timeline Retrieval and Timeline Refinement (Phase II)
Drawing motivation from Figure 1, [Uncaptioned image]Theanine augments response generation (RG) with timelines of relevant events to address the information loss caused by conventional memory management (Xu et al., 2022a; Bae et al., 2022). With $G^{t+1}$, Theanine performs these steps for RG in session $t+1$:
@@
During the conversation, using the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ of $n$ utterances $u$ as query, we first retrieve top-$k$ memories $M_{re}=\{m_{re1},...,m_{rek}\}$.
Phase II-1: Retrieving raw memory timelines.
After that, we wish to also access memories that are centered around relevant events. Formally, given an $m_{re}\in M_{re}$, we further retrieve the connected component $C_{re}\subset G^{t+1}$ that contains $m_{re}$.
@@
where $\Theta$ indicates “the oldest memory in”
@@
Next, starting from $m_{start}$, we untangle the memories by tracing through the future direction and extract every possible linear graph containing $m_{re}$ (two in Figure 4) from $C_{re}$, until reaching an endpoint $\tau[-1]$ with an out-degree of 0 (i.e., $deg^{+}(\tau[-1])=0$, which means no directed edge goes out from it). Each of them is considered a raw memory timeline $\tau$, demonstrating a version of the evolution of $m_{re}$ and its relevant events:
@@
We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^5] Based on our experiments, we empirically set $n$ to 1. Repeating Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
@@
Although we have constructed the memory graph using temporal and commonsense relations to improve its informativeness, directly applying the retrieved timelines for response generation can be suboptimal (RQ3, Section 4), because graph construction does not take the current conversation into consideration, i.e., they are constructed off-line.
In this phase, [Uncaptioned image]Theanine tackles such a discrepancy between off-line memory construction and online deployment (i.e., an ongoing current conversation) via a context-aware timeline refinement. Motivated how LLMs can self-refine their previous generation (Madaan et al., 2024). We leverage LLMs to refine raw timelines into a rich resource of information crafted for the current conversation, by removing redundant information or highlighting information that can come in handy. Formally, given the current dialogue $\mathcal{D}$ and the retrieved raw memory timelines $\mathbb{T}$, the LLM tailors all $\tau\in\mathbb{T}$ into refined timelines $\mathbb{T}_{\Phi}$:
@@ 2.3 Timeline-augmented Response Generation (Phase III)
Finally, we move on to response generation. Here, [Uncaptioned image]Theanine leverages LLMs’ chain-of-thought (CoT) reasoning to discover user intention from the current conversation (Chae et al., 2023) and selectively extract useful information provided in the refined timelines. Formally, given the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ and refined timelines $\mathbb{T}_{\Phi}$, the LLM sequentially generates a CoT rationale $\mathcal{R}$ demonstrating how it comprehends $\mathcal{D}$ and $\mathbb{T}_{\Phi}$, and the next response $\bar{u}_{t+1}$:
¯$\bar{\mathcal{R}}=\underset{\mathcal{R}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{R}|\mathcal{D},\mathbb{T}_{\Phi})$ (12)
$\Rightarrow\bar{u}_{n+1}=\underset{u_{n+1}}{\text{argmax}}\,P_{\text{LLM}}({u}_{n+1}|\mathcal{D},\mathbb{T}_{\Phi},\bar{\mathcal{R}})$ (13)
where $\Rightarrow$ denotes sequential text generation.
@@ 3.1 Datasets of Long-term Conversations
There are few datasets of long-term conversations. First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon the conversations from Persona-Chat (Zhang et al., 2018) by extending their follow-up conversations throughout multiple sessions. Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean, respectively. Recently, Jang et al. (2023) release a new multi-session datasets, Conversation Chronicles (CC). Different from MSC, speakers in CC are augmented with diverse relationships, such as “employee and boss”. Apart from the above open-domain datasets, there is also Psychological QA, [^6] https://www.xinli001.com/ which addresses long-term conversation in Mandarin regarding clinical scenarios.
In this work, we adopt MSC and CC for evaluation. Since we focus on English conversations, we do not include DuleMon, CareCall, and Psychological QA. We leave such multilingual settings and clinical scenarios to future work.
Datasets: Multi-session Chat (MSC) & Conversation Chronicles (CC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 71% 82% 77% 73% 81% 77% 72% 79% 77% 72% 81% 77%
All Memories and $\mathcal{D}$ 69% 64% 66% 69% 69% 68% 70% 68% 68% 69% 67% 67%
     + Memory Update 56% 57% 60% 56% 56% 58% 56% 58% 60% 56% 57% 60%
Memory Retrieval 67% 62% 68% 72% 67% 68% 72% 65% 68% 70% 65% 68%
     + Memory Update 57% 56% 59% 58% 57% 61% 55% 58% 59% 57% 57% 60%
RSum-LLM 38% 61% 69% 35% 57% 67% 37% 57% 65% 37% 58% 67%
MemoChat 60% 74% 71% 60% 69% 72% 60% 72% 71% 60% 72% 71%
COMEDY 66% 54% 78% 64% 54% 77% 62% 53% 76% 64% 53% 77%
Golden Responses 63% 76% 61% 63% 73% 61% 62% 72% 61% 63% 73% 61%
Table 1: Win rates of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Spe = specificity; Mem = memorability; Cst = consistency. Color green denotes metrics regarding the reference of memories. Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.

### v2

# Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation

@@ Abstract
To achieve lifelong human-agent interaction, dialogue agents need to constantly memorize perceived information and properly retrieve it for response generation (RG). While prior work focuses on getting rid of outdated memories to improve retrieval quality, we argue that such memories provide rich, important contextual cues for RG (e.g., changes in user behaviors) in long-term conversations. We present [Uncaptioned image]Theanine, a framework for LLM-based lifelong dialogue agents. Theanine discards memory removal and manages large-scale memories by linking them based on their temporal and cause-effect relation. Enabled by this linking structure, Theanine augments RG with memory timelines - series of memories representing the evolution or causality of relevant past events. Along with Theanine, we introduce TeaFarm, a counterfactual-driven evaluation scheme, addressing the limitation of G-Eval and human efforts in measuring memory-augmented dialogue agents. A supplementary video for Theanine and data for TeaFarm are at https://huggingface.co/spaces/ResearcherScholar/Theanine.
Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation
Kai Tzu-iunn Ong1thanks: KT Ong and N Kim are the co-first authors.    Namyoung Kim1    Minju Gwak1    Hyungjoo Chae1 Taeyoon Kwon1    Yohan Jo2    Seung-won Hwang2    Dongha Lee1    Jinyoung Yeo1 1Yonsei University, 2Seoul National University
@@ 1 Introduction
Autonomous agents based on large language models (LLMs) have made great progress in various domains, including response generation (Zhou et al., 2023; Kwon et al., 2024; Tseng et al., 2024), where agents ought to constantly keep track of both old and newly introduced information shared with users throughout their service lives (Irfan et al., 2024) and converse accordingly. To facilitate such lifelong interaction, studies have proposed to improve dialogue agents’ ability to memorize and correctly recall past information (e.g., discussed topics) in long-term, multi-session conversations.
Figure 1: Empirical examples of failed responses due to (a) absence of an important past event (“afraid of cruise ships”) on the timeline and (b) bias to the latest input. (c) is a response augmented with the memory timeline.
@@
A representative method is to compress past conversations into summarized memories and retrieve them to augment response generation (RG) in later encounters (Xu et al., 2022a; Lu et al., 2023). However, the growing span of memories can hinder retrieval quality as conversations accumulate. Although it, to some extent, can be solved by updating old memories (Bae et al., 2022; Zhong et al., 2024), this common practice may cause severe information loss. As shown in Figure 1 (a), an earlier memory on the timeline, an important persona (“afraid of ships”), is removed during memory update and causes improper RG. While using the large context windows of recent LLMs to process all dialogue history/memories is an option to prevent such information loss, [^1] For instance, GPT-4o and Llama 3.1 have context windows of 128K tokens (OpenAI, 2024a; MetaAI, 2024). this often leads to biased attention toward the latest user input (Figure 1 (b)), ignoring relevant contexts from the past (Liu et al., 2024). These findings highlight two main challenges towards lifelong dialogue agents - (i) Memory construction: how to effectively store large-scale past interactions without removing old memories? (ii) Response generation: within the growing memory span, how to identify relevant contextual cues for generating proper responses?
Motivated by these, we propose to address the above two challenges separately yet complementarily, by (i) discarding memory update to avoid information loss, and preserving relevant memories on the timeline in a linked structure; and (ii) retrieving the timeline as a whole to better catch relevant memories within the growing search span. We present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019). a framework for facilitating lifelong dialogue agents.
Starting from memory construction (Phase I), instead of stacking raw memory sentences as-is (Xu et al., 2022a), which may affect memory retrieval and also response quality due to the unstructured format of information (Mousavi et al., 2023; Chen et al., 2023), Theanine stores memories in a directed graph. In this graph, inspired by how humans naturally link new memories to existing ones of relevant events based on their relation (Bartlett, 1995), memories are linked using their temporal and cause-effect commonsense relations (Hwang et al., 2021). Supported by such linking structure, in memory retrieval for RG (Phase II-1), we go beyond conventional top-$k$ retrieval and further obtain the complete timelines to avoid missing out on important memories that have low textual overlap with current conversation (Tao et al., 2023). Lastly, to tackle the discrepancy between off-line memory construction and online deployment, Theanine uses an LLM to refine retrieved timelines (Phase II-2) based on current conversation (Madaan et al., 2024), such that they provide tailored information (Chae et al., 2023) for generating responses (Phase III). Our contributions are two-fold:
  • To achieve lifelong dialogue agents, we present [Uncaptioned image]Theanine, an LLM-based framework with a relation-aware memory graph and timeline-augmented response generation for long-term conversations. Theanine outperforms baselines across automatic, LLM-based, and human evaluations of RG. Also, we confirm that Theanine leads to better retrieval quality, and its intermediate procedures align with human preference. To our knowledge, we are the first to model the concept of timelines in memory management and RG.
@@
  • The lack of golden mapping between conversations and reference memories poses a challenge in assessing memory-augmented agents. We present TeaFarm, a counterfactual-driven pipeline evaluating agent performance in referencing the past without human intervention.
@@ 2 Methodologies
We present [Uncaptioned image]Theanine, a framework for lifelong dialogue agents, inspired by how humans store and retrieve memories for conversations (Figure 2):
@@ 2.1 Memory Graph Construction (Phase I)
To manage large-scale memories and facilitate structured information for RG (Mousavi et al., 2023; Chen et al., 2023), we approach memory management using a memory graph $G$:
@@
Humans, on the other hand, interpret events by considering the relation between them, such as “how does an event affect the other?” or “why did this person make that change?”. Therefore, we adopt a relation-aware memory linking, where an edge between two memories is encoded with their cause-effect commonsense relation $r\in R$, along w/ the temporal order. In practice, we adopt the commonly used relations defined by Hwang et al. (2021), including HinderedBy, Cause, Want, and 4 more (Appendix B.1).
@@
where $\Upsilon(\cdot,m_{new})\in R$ indicates that the given memory is assigned with an $r\in R$ with $m_{new}$, [^4] Limited by retrievers, an $m\in M_{a}$ may not have a relation with $m_{new}$. We thus allow the LLM to output “None”. and such assigned memories are defined as $M_{a}^{*}$.
We then proceed to link $m_{new}$ to the graph. We first locate every connected component $C_{i}\subset G^{t}$ that contains at least one $m\in M_{a}^{*}$, as shown in Figure 3 (a) and (b):
where $\mathbb{C}$ is the collection of those $C$ and $\mathtt{V}(\cdot)$ represents “vertices in”. Then, we link $m_{new}$ to the most recent [^5] Simply linking $m_{new}$ to all $m\in M_{a}^{*}$ costing 25% more API cost for linking without leading to better response. $m\in M_{a}^{*}$ in each $C_{i}\subset\mathbb{C}$ (Figure 3 (c)). Memories $M_{linked}$ that are linked to $m_{new}$ is defined as follows:
where $\Omega(\cdot)$ indicates “the most recent memory in”.
@@
Linking all memories from session $t$ to $G^{t}$, we then obtain a new memory graph $G^{t+1}$. The pseudo algorithm for Phase I is in Algorithm 1.
@@ 2.2 Timeline Retrieval and Timeline Refinement (Phase II)
Thanks to the constructed memory graph, Theanine can proceed to augment RG with timelines of relevant events, addressing the information loss in conventional memory management (Figure 1). With $G^{t+1}$, Theanine performs the following steps for RG in session $t+1$:
@@
During the conversation, using the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ of $n$ utterances $u$ as query, we retrieve top-$k$ memories $M_{re}=\{m_{re1},...,m_{rek}\}$.
Phase II-1: Retrieving and untangling raw memory timelines.
We wish to also access memories centered around $M_{re}$. Formally, given $m_{re}\in M_{re}$, we further collect the connected component $C_{re}\subset G^{t+1}$ that contains $m_{re}$ via the linked structure.
@@
where $\Theta$ indicates “the oldest memory in”.
@@
Next, starting from $m_{start}$, we untangle memories by tracing through future direction and extract every possible linear graph containing $m_{re}$ (two in Figure 4) from $C_{re}$, until reaching an endpoint $\tau[-1]$ with an out-degree of 0 (i.e., $deg^{+}(\tau[-1])=0$), which means no directed edge goes out from it). Each of them is considered a raw memory timeline $\tau$, demonstrating a version of the evolution of $m_{re}$ and its relevant events:
@@
We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^6] Based on our experiments, we empirically set $n$ to 1. Repeating777“Repeating” is used to explain the algorithm from the perspective of one $m_{re}$. In practice, $M_{re}$ are processed together, although processing them 1-by-1 yields the same result. Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
@@
Although we have constructed the memory graph using temporal and commonsense relations to improve its informativeness, directly applying retrieved timelines for RG can be suboptimal (RQ3, Section 4), because graph construction does not take current conversation into consideration, i.e., they are constructed off-line.
In this phase, Theanine tackles such a discrepancy between off-line memory construction and online deployment (i.e., ongoing conversation) via a context-aware timeline refinement. Motivated by how LLMs can refine their previous generation (Madaan et al., 2024). We leverage LLMs to refine raw timelines into a rich resource of information crafted for the current conversation, by removing redundant information or highlighting information that can come in handy. Formally, given the current dialogue $\mathcal{D}$ and retrieved raw timelines $\mathbb{T}$, an LLM tailors $\tau\in\mathbb{T}$ into refined timelines $\mathbb{T}_{\Phi}$:
@@ 2.3 Timeline-augmented Response Generation (Phase III)
Now, Theanine utilizes the refined timelines for RG. Formally, given $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ and $\mathbb{T}_{\Phi}$, an LLM generates a next response $\bar{u}_{t+1}$:
$\bar{u}_{n+1}=\underset{u_{n+1}}{\text{argmax}}\,P_{\text{LLM}}({u}_{n+1}|\mathcal{D},\mathbb{T}_{\Phi})$ (12)
@@ 3 Experimental Setups
Datasets: Multi-session Chat (MSC) Conversation Chronicles (CC)
Methods / Metrics Bleu-4 Rouge-L Mauve BertScore Bleu-4 Rouge-L Mauve BertScore
All Dialogue History 1.65 14.89 9.06 86.28 4.90 21.56 26.47 88.13
All Memories & Current Context $\mathcal{D}$ 1.56 14.89 10.62 86.23 4.41 20.06 38.16 88.02
+ Memory Update (Bae et al., 2022) 1.55 14.77 9.28 86.20 4.34 20.34 34.84 88.03
Memory Retrieval (Xu et al., 2022a) 1.92 15.49 11.16 86.47 4.93 20.63 33.06 88.07
+ Memory Update (Bae et al., 2022) 1.67 15.30 13.71 86.39 4.46 20.19 34.28 88.02
Rsum-LLM (Wang et al., 2023) 0.75 11.53 2.45 84.91 0.98 11.42 2.28 85.59
MemoChat (Lu et al., 2023) 1.42 13.51 7.72 85.96 2.31 15.87 15.12 87.08
COMEDY (Chen et al., 2024b) 1.06 12.79 7.27 85.29 1.70 13.57 1.95 85.90
[Uncaptioned image]Theanine (Ours) 1.80 15.37 18.62 86.70 6.85 22.68 64.41 88.58
Table 1: Automatic evaluation of response quality (average of sessions). Session-specific results are in Appendix F.
@@ 3.1 Datasets of Long-term Conversations
There are few datasets of long-term, multi-session conversations. First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon Persona-Chat (Zhang et al., 2018) by extending its conversations to multiple (five) sessions. Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean. Recently, Jang et al. (2023) release a new dataset, Conversation Chronicles (CC). Different from MSC, speakers in CC are augmented with relationships, such as “employee and boss”. Apart from the above open-domain datasets, there is also Psychological QA, [^8] https://www.xinli001.com/ which addresses long-term conversations in Mandarin regarding clinical scenarios.
We use MSC and CC for evaluation. We focus on English conversations, so we do not adopt DuleMon, CareCall, and Psychological QA and leave multilingual and clinical settings to future work.